# Instruction Fine-Tuning: Does Prompt Loss Matter?

### v2

@@ 7 Conclusion
In this study, we explored the downstream effects of prompt loss weight (PLW) for LLM supervised instruction fine-tuning (SIFT).
We showed that performance is unaffected for the two long-completion datasets (AlpacaData, the original Alpaca dataset, and AlpacaDataCleaned, an updated version of AlpacaData) while performance had a negative quadratic relationship with prompt loss weight on our short-completion dataset (with a completion-prompt ratio $R_{g}\approx 0.08$).
Interestingly, the highest scores on the ARC Challenge, PIQA, and WinoGrande benchmarks were all from models trained on short-completion dataset. This suggests that given a good choice of prompt loss weight, long-completion training data may not be necessary for instruction fine-tuning LLMs for language understanding and short generation tasks. Conversely, using long-completion data should be more robust as prompt loss weight and masking parameters can be ignored.
We further analyzed the AlpacaDataShort models and found that low non-zero PLW discourages deviation from the PTLM weights and high PLW reduces overfitting and encourages longer generation. We suggest that these two factors are responsible for the negative quadratic relationship between performance and PLW.
Finally, we estimated optimal PLWs over a range of completion-prompt length ratios.
@@
    We used a fixed seed $=42$ for all experiments. Since we used pre-trained models and no layers were freshly initialized, there was little variance in initial experiments. Therefore, we decided to limit runs to a single seed.
@@
    Suggested values for PLW from section 6 are based on the included experiments. Best PLW values when fine-tuning different models or using different datasets or training regimes may vary from the relationships shown here, though we are still confident that performance will not vary significantly by PLW for long-completion data.
@@
    LLM-as-evaluator approaches like PandaLM and Alpaca Eval 1 are still relatively new, and these approaches are still being actively developed. We chose to use Mixtral as an auto-evaluator for Alpaca Eval. Although this is not the recommended auto-evaluator, we felt its performance on the human-labeled test set was good enough to justify its use.
@@
    A more confident conclusion about the causative mechanism could not be made. As with most techniques in machine learning, the suggested causative relationship should be validated through numerous experiments and projects. Unfortunately, since we expect most instruction fine-tuning datasets to be long-completion datasets, further research into PLW will likely be limited.
@@
We do acknowledge that standard deep learning training methods have a high carbon footprint, and we performed a total of 100 fine-tuning training runs. Model outputs cannot be predicted in advance, and, while we release our model weights in the spirit of transparency and collaboration, models may hallucinate or produce offensive output. Our synthetic dataset was generated from another publicly released dataset and we did not perform additional content filtering. A warning about both of these issues will be released along with the model and dataset.
@@ Acknowledgements
This research was supported by Gwangju City’s Artificial Intelligence Industry Cluster Agency (AICA) and by the Korean Ministry of Trade, Energy, and Industry (MOTIE) under the KoSIGN project [Grant Number: 20014406].
This research was possible thanks to the support of my coworkers at EQ4ALL.
Appendix A Benchmark Performance Figures
This appendix contains individual benchmark performance plots and additional discussion of performance trends and results.
Overall, models trained on AlpacaDataCleaned performed best on the selected benchmarks, with models trained on AlpacaDataShort meeting or outperforming these models on numerous benchmarks for optimal selection of PLW. The general PTLM trend was for LLaMA 2 models to outperform LLaMA 1 models, again with some exceptions. In general, models trained on AlpacaData and AlpacaDataCleaned did not show consistent relationships with PLW on the tested benchmark.
@@ Appendix A Benchmark Performance Figures
The rest of this section focuses on AlpacaDatashort models. We divide the benchmarks into three groups based on the relationship between AlpacaDataShort model performance and PLW.
Group I benchmarks showed a negative quadratic relationship with the transformed PLW variable wp: ARC Challenge, PIQA, TruthfulQA-Gen, and WinoGrande. See figure 4 for results.
Group II benchmarks showed an increasing relationship between performance and wp on AlpacaDataShort: TruthfulQA-MC2, Alpaca Eval 1 (Mixtral), and PandaLM. Interestingly, Alpaca Eval 1 (Mixtral) and PandaLM are both long generation benchmarks, but the generation version of the TruthfulQA benchmark is in the first group while the multiple choice benchmark is in this group.
Group III are high noise benchmarks, limited to the six WMT translation tasks. See figure 6 for individual plots. In general WMT tasks showed very small variance, at most varying across only a few percentage points. To reduce the noise, these benchmarks were calculated over both the validation and test sets, but noise still precludes making clear conclusions about performance on any benchmark. Of note, combining benchmarks that translate from English produces a trend line suggestive of a positive relationship with wp. See figure 5(a) for an example. Figure 5(b) portrays into English translation benchmarks, but no trends could be identified.
@@
Appendix B Recommended Prompt Loss Weight
In this section, we present additional figures and tables related to optimal prompt loss weights (PLWs) as predicted by fitted generalized additive models (GAMs).
The optimal PLW predictions from the paper body were based on all benchmarks. Here we present three different relative aggregates: all benchmarks (labeled “All”), multiple choice benchmarks (labeled “MC”), and non-translation generation benchmarks (labeled “Gen”). We found that the translation benchmarks contributed little to the predictive power of the fitted GAMs. In an attempt to predict PLW values in a more relevant context, we did not include the translation benchmarks in the “Gen” model. However, they are still included in the “All” model to capture a wider range of tasks.
See figures 7,8,9 for contour plots for the “All”, “MC”, and “Gen” benchmarks, respectively. The “All” contour plot is identical to the contour plot in the main body.
See table 4 for a list of optimal PLWs over a range of completion-prompt ratios. Note that the predicted optimal PLWs confirm the conclusions from our regression analysis in section 5.2. Namely that PLW is important for short-completion data, but can be ignored for long-completion data.
@@ Appendix B Recommended Prompt Loss Weight
Figure 7: “All” Ratio-PLW Interaction for all benchmarks.
Figure 8: “MC” Ratio-PLW Interaction for multiple choice benchmarks.
Figure 9: “Gen” Ratio-PLW Interaction for non-translation generation benchmarks.
Appendix C Reproducibility
This section provides technical details on all experiments and benchmarks for transparency and to encourage reproduction of results. To help with reproducibility, we will also upload the fine-tuned models, test generation outputs, and the AlpacaDataShort dataset to the HuggingFace Hub.
C.1 Model Fine-Tuning
@@
AlpacaData is available from the Stanford Alpaca repository. AlpacaDataCleaned can be found at www.github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”. As noted above, we will release AlpacaDataShort in the near future to the HuggingFace Hub.
C.2 Model Evaluation
@@
TruthfulQA-Gen and all translation tasks were evaluated using a more recent commit at www.github.com/EleutherAI/lm-evaluation-harness/tree/b93c3bc. We modified the translation tasks at this commit to include an appropriate prompt to support zero-shot translation.
@@ C.4 Causal Mechanism
C.5 Predictive Model
We fit several generalized additive models (GAMs) using the mgcv library, version 1.9-1 and the same version of R as above, version 4.3.0.
Appendix D Artifact Licensing

### v3

@@ 7 Conclusion
In this study, we explored the effects of prompt loss weight (PLW) on LLM supervised instruction fine-tuning (SIFT).
We found that PLW has a statistically significant effect on learning for our short-completion dataset, and proper tuning of PLW allowed short-completion-trained models to outperform long-completion-trained models on short-generation benchmarks. We showed that the causal mechanism was due to a balance between two different regularizing effects and not due to increased training stability as is commonly attributed. We showed that the measured relationship extends to additional SIFT datasets and that the effects could not be sufficiently emulated with other regularizers.
Based on the above conclusions, we assert the following two points.
  1. 1.
    Since models fine-tuned on short-completion datasets and with properly tuned PLWs outperformed all other models on short-generation benchmark tasks, we conclude that PLW is critical for effectively fine-tuning for downstream short-generation tasks.
  2. 2.
    Given the importance of PLW and given that many SIFT datasets and almost all natural language understanding (NLU) datasets are short-completion datasets, we warn SFT API providers about the need for a PLW parameter to adequately cover a full range of modeling applications.
@@
    Since we used pre-trained models and no layers were freshly initialized, there was little variance in initial experiments. We therefore limited runs to a single seed of $42$.
@@
    Suggested values for PLW from section D are based on the included experiments. Best PLW values when fine-tuning different models or using different datasets or training regimes may vary from the relationships shown here, though we are still confident that performance will not vary significantly by PLW for long-completion data.
@@
    The focus of our research was on how PLW affected fine-tuning based on the completion-prompt ratio of the training dataset. However, the absolute length and size of the dataset will likely play a role in learning dynamics. It would be good to include that perspective in future research on token loss weights.
@@
    LLM-as-evaluator approaches like PandaLM and AlpacaEval 1 are still relatively new, and these approaches are being actively developed. We chose to use Mixtral 8x7B as an auto-evaluator for AlpacaEval 1 due to budget limitations. While we cite high human evaluation correlation with Mixtral and justify this decision in section 4.4, using the default auto-evaluator would be beneficial for better comparison with other research.
  6. 6.
    While we define short- and long-completion data as have a completion-prompt ratio $R_{g}$ lower and greater than $1$, respectively, we do not provide justification for using $1$ as the threshold. Choosing a meaningful reference would be helpful to future research.
@@
We acknowledge that standard deep learning training methods have a high carbon footprint, and we performed over 200 fine-tuning training runs. Model outputs cannot be predicted in advance, and, while we release our model weights in the spirit of transparency and collaboration, models may hallucinate or produce offensive output. Additionally, our shortened datasets were generated from publicly released data, and we did not perform additional content filtering. A warning about both of these issues will be released along with the models and datasets.
@@ Acknowledgements
This work was supported by the Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City (No. BA00000797, LLM-based sign language translation for weather forecasts).
This research was possible thanks to the support of our coworkers at EQ4ALL.
Appendices
Appendix A Prompt Inversion
Figure 4: Examples of modifying prompt-completion ratios using prompt inversion, best viewed in color. To prompt-invert instances, we re-frame the prompt-completion task as an original-prompt-prediction task. I.e., we teach the model to predict the original instruction given an example completion and optional input. In the first example above, prompt inversion changes the instance’s word-based completion-prompt ratio $R_{g}$ from $34/(7+0)=4.857$ to $7/(9+34)=0.163$.
In order to experiment with short-completion fine-tuning datasets while also preserving a similar dataset to use as a reference, we propose using prompt inversion to rotate the instruction, input, and output fields of instances. Prompt inversion modifies an instance to use the original instruction as the new completion text and the original output as the new input text. The model is then given the following instruction:
Predict the prompt that generated the following AI output.
if the input field is empty and
Predict the prompt that generated the below AI output given the following context. Context: <original-input>
if there is an input field. See figure 4 for a visualization of this process.
To generate short versions of our test instruction datasets, we use prompt inversion to modify every instance with a completion-prompt length ratio $R_{g}>1$, based on the tokenized lengths of each field and using prompt to mean the concatenation of the instruction and input fields as explained in section 2.1.
Synthesis of the original input and output texts to predict the original instruction requires both language understanding and reasoning. It also does not train the model for long-generation (since new completions will generally be short) and thus does not prepare the model for train-of-thought generation that long-generation tasks intrinsically learn. Therefore, we propose that prompt-inverted instances are best suited for learning natural language understanding (NLU)-related tasks rather than general generation tasks.
Appendix B Main Experiment Benchmarks
This section presents additional qualitative analysis of benchmark performance and score visualizations for each benchmark.
@@ Appendix B Main Experiment Benchmarks
For both the simple aggregate and the relative aggregate, models trained on AlpacaDataShort showed a visual relationship with wp. Based on this visual relationship, we divide benchmarks into three groups.
The first group showed a negative quadratic relationship with wp, with performance exceeding that of AlpacaDataCleaned models. This group consists of ARC Challenge, PIQA, TruthfulQA-Gen, and WinoGrande benchmarks, and optimal PLW values for these four benchmarks vary from PLW $=0.01$ to PLW $=0.1$. See figure 5 for individual benchmark visualizations.
The second group of benchmarks showed steadily increasing performance as wp increased, before leveling off to maximum values near wp$=1$. This group is TruthfulQA-MC2, AlpacaEval 1, and PandaLM. It is surprising that TruthfulQA-MC2 shows a relationshp more similar to the long-generation benchmarks and TruthfulQA-Gen resembles the other multi-choice benchmarks. See figure 6 for individual benchmark visualizations.
[…]
The third group consists of the six translation benchmarks and showed unclear correlation between performance and wp. Though aggregating benchmarks into “to English” and “from English” subgroups creates visualizations suggestive of a relationship, benchmarks from this group showed relatively more noise than the other benchmarks. To reduce score noise, translation benchmarks were evaluated on the combined validation and test data splits, but there was still significant noise in the results. See figure 7 for individual benchmark visualizations.
@@
The performance difference across different wp values for the two long-generation benchmarks was around twenty percentage points, in stark contrast to the less than two percentage point change for short-generation and multiple choice benchmarks. This suggests that PLW plays an important role in the ability to generate high quality text, and the optimal PLW for short-generation and long-generation benchmarks is clearly different. Also note that performance of LLaMA 2 models was in general higher than that of LLaMA 1 models and performance of AlpacaDataCleaned models were higher than that of AlpacaData models, validating the improvements of LLaMA 2 and AlpacaDataCleaned over their predecessors.
@@
Appendix C Visualizations for Supplemental Experiments
This appendix contains visualizations and additional details about the two supplemental experiments from section 6.
(a) Simple Aggregate
(b) Relative Aggregate
(c) Relative Aggregate (Group I)
(d) Relative Aggregate (Group II)
Figure 8: Comparison of PLW with other regularization techniques (calculated for PLW $=0$ and PLW $=1$). (a) The simple aggregate. (b) The combined relative aggregate shows that models fine-tuned with fractional PLW on AlpacaDataShort outperformed models fine-tuned with alternative regularizations. (c) Fractional PLW performance is most extreme for multiple choice and short-generation benchmarks (group I). (d) Performance of fractional PLW models on group II benchmarks is less pronounced, with PLW-optimized models performing slightly worse than several other alternative metrics.
C.1 Regularization Comparison
For the first supplemental experiment, we wanted to investigate if PLW is necessary for fine-tuning on short-completion data or if another regularization technique could yield the same benefits. As explained in section 6.1, we chose to examine four types of regularization in addition to PLW, intentionally not evaluating KL divergence-based regularization due to the difficulty of applying it to LLM SIFT.
Several aggregate scores for regularizations with example parameters are presented in figure 8, and best scores for each type of regularization is presented in table 4 in the main paper. Visualization of relative aggregate scores revealed that models fine-tuned with fractional PLW generated high scores on multiple-choice and short-generation benchmarks while long-generation benchmarks (AlpacaEval 1 and PandaLM) actually benefitted the most from alternative regularization methods. However, the effect on the multiple choice and short generation benchmarks was relatively strong, and the combined relative aggregate also showed maximal values for models fine-tuned with fractional PLW.
Interestingly, most regularization methods performed better when coupled with PLW$=1$ than with PLW$=0$ except for label smoothing which performed marginally worse at PLW$=1$ than at PLW$=0$.
C.2 Dataset Comparison
(a) Original Datasets
(b) Short Dataset Variants
(c) UltraFeedbackBinarizedClean
(d) UltraFeedbackBinarizedShort
(e) DatabricksDolly
(f) DatabricksDollyShort
Figure 9: Relative aggregate scores for models fine-tuned on alternative instruction datasets.
In the second supplemental experiment, we wanted to explore if the relationship between PLW and model performance on downstream tasks measured for AlpacaDataShort models existed for other fine-tuning datasets as well. See additional visualizations in figure 9.
Appendix D Optimal Prompt Loss Weight
(a) “All” Prediction
(b) “MC” Prediction
(c) “Gen” Prediction
Figure 10: Best viewed digitally for improved resolution.
In the main regression experiment, we showed that PLW is an important hyperparameter when fine-tuning on short-completion data but is effectively irrelevant when using long-completion data. In this appendix, we present several models for predicting an optimal PLW given a dataset’s $R_{g}$. These models are based on the AlpacaData dataset, AlpacaDataCleaned dataset, and several modified versions of AlpacaDataCleaned and therefore should be seen as an exercise rather than an authoritative reference on optimal PLW weights.
We first repeated our SIFT experiments on two additional datasets: AlpacaDataMedium and AlpacaDataTiny to increase coverage of the parameter space. AlpacaDataMedium and AlpacaDataTiny have $R_{g}$ values of 1.0 and 0.042, respectively, and were generated using prompt inversion (see Appendix A) but selecting instances to modify in order to approach target $R_{g}$ values.
@@ Appendix D Optimal Prompt Loss Weight
Using the fitted w-r interaction, we then estimated optimal PLW value for a given completion-prompt ratio $R_{g}$. See figure 10(a) for a visualization of a GAM fitted on all benchmark tasks.
Roughly, the fitted interaction term recommended using PLW $=0.155$ for small completion-prompt length ratios ($R_{g}\leq 1$) and up to PLW $=0.278$ for a $R_{g}=1.5$ for optimal performance across all tasks. This prediction is close to our regression-predicted value of 0.242. The interaction term also confirms our observations that PLW is less important for data with relatively long completions.
Since the relationship between PLW and benchmark performance depends heavily on the type of benchmark task, we also fit GAMs for an aggregate of multiple choice benchmark scores (labeled “MC”) and generation benchmark scores (labeled “Gen”). We found that the translation benchmarks contributed little to the predictive power of the fitted GAMs and while their scores are included in the “All” GAM, we did not include them when fitting the “Gen” GAM. See figures 10(b) and 10(c) for contour plots for the “MC” and “Gen” benchmarks, respectively.
Also see table 5 for a list of GAM-based optimal PLWs over a range of completion-prompt ratios. Again, predicted optimal PLWs confirmed the conclusions from our regression analysis in section 5.2.
Appendix E Reproducibility
This section provides technical details on all experiments and benchmarks for transparency and to encourage reproduction of results. To help with reproducibility, we will also upload the fine-tuned models, test generation outputs, and our modified datasets to xx.yy.zz.
E.1 Model Fine-Tuning
@@
AlpacaData is available from the Stanford Alpaca repository. AlpacaDataCleaned can be found at www.github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”. As noted above, we released AlpacaDataShort at xx.yy.zz.
E.2 Model Evaluation
@@
TruthfulQA-Gen and all translation tasks were evaluated using a more recent commit at www.github.com/EleutherAI/lm-evaluation-harness/tree/b93c3bc. We modified the translation tasks at this commit to include an appropriate prompt to support zero-shot translation. These changes can be seen at xx.yy.zz.
@@ E.4 Causal Mechanism
E.5 Supplemental Experiments
The first supplemental experiment used the same experimental setup and data from the main experiment. Of the tested regularization methods, we used the weight decay implementation in PyTorch’s AdamW optimizer, attention dropout implemented by the LlamaAttention module from Transformers, and label smoothing supported by the Trainer class from Transformers. We manually implemented regularization based on the Minkowski distance by calculating the mean of the $p=1$ Minkowski distance between each pair of weight tensors from the PTLM and the trained model.
The second experiment introduced two new datasets: UltraFeedbackBinarizedCleaned and DatabricksDolly. UltraFeedbackBinarizedCleaned can be found at https://huggingface.co/datasets/allenai/ultrafeedback_binarized_cleaned/tree/f304ce5. And DatabricksDolly can be found at https://huggingface.co/datasets/databricks/databricks-dolly-15k/tree/bdd27f4.
E.6 Predictive Model
We fit several generalized additive models (GAMs) in Appendix D using the mgcv library, version 1.9-1 and the same version of R as above, version 4.3.0.
Appendix F Artifact Licensing
@@
AlpacaDataCleaned Apache 2.0 Model Training
Ultrafeedback Binarized Cleaned MIT Model Training
databricks-dolly-15k CC BY-SA 3.0 Model Training
@@
LLaMA 3 LLaMA 3 Community License Pre-trained model weights